Consider the UNILUS economy in the long-run described by the following equations:

Y = C + I + G, Y = 5,000, G = 1,000, T = 1,000, C = 250 + 0.75(Y − T), I = 1,000 − 50r. C=100+b(Y-50-0.25Y);

Compute private saving, public saving, and national saving.

(5 Marks)

Find the equilibrium interest rate. (5 Marks)

Now suppose that G rises to 1,250. Compute private saving, public saving, and national savings. (5 Marks)

Find the new equilibrium i


1
Expert's answer
2021-09-13T12:41:34-0400

"Part\space (a)"


"Y=C+I+G"


 "C=250+0.75(5000-1000)\n=250+3000=3250"

"5000=3250+1000-50r+1000"

"5000=5250-50r"

"50r=250"

"r=250\div50=5"

"I=1000-(50\times5)=1000-250=750"


Private savings.

"S=Y-C-G=I=5000-3250-1000=750"


Public savings

"Sg=T-G=1000-1000=0(Balanced\space budget)"

 "Sp=-250+(1-0.75)(Y-T)=-250+0.25(5000-1000)=-250+1000=750"


National savings

"S=Sp=Sg=750+0=750 (National)"



"Part (b)"


"S=I"

"750=1000-50r"

 

"r=5\%"


"Part (c)"


"Y=C+I+G"


"C=3250"


"5000=3250+1000-50r+1250"


 "5000=5500-50r"


 "50r=500"


"r=500\div50 =10"


 "I=1000-(50\times10)=1000-500=500"


 "S=Y-C-G=I=5000-3250-1250=500"

Public savings

 "Sg=T-G=1000-1250=-250 \space(Budget\space deficit)"

"Public"

Private savings

"Sp=Y-T-C=5000-1000-3250 =750"

"Sp=750\space (Private)"


 National savings

"S=Sp+Sg=750-250=500\space (National)"


 "Part\space (d)"

New equilibrium interest


"Y=C+I+G"


"5000=3250+1000\u221250r+1250"

"5000=5500-50r"

 "50r=500"

"r=500\div50 =10"


"r=10\%"